### v1

# On the interplay between the BFKL resummation and high-energy factorization in Mueller-Navelet dijet production

Abstract
We study the interplay between the high-energy resummation based on the BFKL equation and initial state radiation effects, taken into account within the framework of the high-energy factorization, in the production of Mueller-Navelet dijets at hadron colliders. We use the solution of the NLO BFKL equation built out of eigenfunctions perturbatively constructed up to NLO to avoid the need for a special renormalization scale setting. We demonstrate that the various data sets from the FNAL Tevatron and CERN LHC can be described in this way and both the initial state radiation effects of the high-energy factorization and the NLL BFKL resummation are crucial for the uniform description of the data across all values of rapidity difference between the jets.
pacs:
12.38.Bx, 12.39.St, 13.85.Tp
The Balitsky-Fadin-Kuraev-Lipatov (BFKL) equation describing the structure of the QCD pomeron has a history of almost $50$ years Lipatov (1976); Kuraev et al. (1976, 1977); Balitsky and Lipatov (1978); Fadin and Lipatov (1998, 2018). It can be written as an evolution equation for the Green’s function $G$ w.r.t. rapidity difference $Y$:
$\frac{\partial}{\partial Y}\ G\left({\bf l}_{T1},{\bf l}_{T2},Y\right)=\int\limits_{{\bf q}_{T}}K\left({\bf l}_{T1},{\bf q}_{T}\right)\,G\left({\bf q}_{T},{\bf l}_{T2},Y\right),$ (1)
where the BFKL kernel $K$ is now known at leading order (LO, ${\cal O}(\alpha_{s})$Balitsky and Lipatov (1978) and next-to-leading order (NLO, ${\cal O}(\alpha_{s}^{2})$Fadin and Lipatov (1998), such that the Green’s function resums the leading logarithmic (LL $\sim\alpha_{s}^{n}Y^{n}$) or next-to-leading logarithmic (NLL $\sim\alpha_{s}^{n+1}Y^{n}$) corrections to the cross section respectively. The calculation of next-to-next-to-leading-order corrections to the kernel is now in progress Fadin and Lipatov (2018); Byrne (2024); Fadin et al. (2023); Byrne et al. (2022); Caola et al. (2022); Falcioni et al. (2022); Del Duca et al. (2022); Abreu et al. (2025).
Many different hard processes were proposed to probe the BFKL dynamics at colliders, see Hentschinski et al. (2023) for an overview. One of the most popular is the inclusive production of a pair of jets with large transverse momenta $|{\bf p}_{T1,2}|\gg\Lambda_{\rm QCD}$ and large invariant mass $M^{2}_{12}\gg|{\bf p}_{T1}||{\bf p}_{T2}|$ widely separated in the rapidity $Y\sim\ln\left(M_{12}^{2}/(|{\bf p}_{T1}||{\bf p}_{T2}|)\right)$, as suggested by Mueller and Navelet (MN) Mueller and Navelet (1987). Several comprehensive studies of the MN dijet production were performed using the hybrid approach of conventional collinear factorization (CF) together with the BFKL resummation in hard scattering coefficient (HSC), see Refs. Andersen et al. (2001); Sabio Vera and Schwennsen (2007); Colferai et al. (2010); Ducloue et al. (2013); Ducloué et al. (2014, 2015); Mueller et al. (2016); Caporale et al. (2013, 2014, 2015); Celiberto and Papa (2022); Egorov and Kim (2023) and references therein. However, there exist well-known difficulties with such hybrid approach associated with the appearance of large logarithmic corrections to impact-factors (IFs) that need to be resummed for a complete description of the MN dijet production.
The high-energy factorization (HEF) formalism Gribov et al. (1983); Collins and Ellis (1991); Catani and Hautmann (1994) was first introduced as a resummation tool for high-order QCD corrections enhanced by logarithms of partonic center-of-mass energy $\ln\left(\hat{s}/\mu^{2}\right)$ within CF. Later, the parton Reggeization approach (PRA) Nefedov et al. (2013); Karpishkov et al. (2017); Nefedov and Saleev (2020) was suggested to define the HEF HSC with off-shell initial states in a gauge-invariant way using high-energy QCD effective field theory (EFT) formulated by L.N. Lipatov Lipatov (1995) and to include the resummation of Sudakov logarithms $\ln\left({\bf q}_{T}^{2}/\mu^{2}\right)$ in the unintegrated PDF (UPDF) Nefedov (2021). Feynman rules of the EFT were derived in Ref. Antonov et al. (2005). The scheme of the NLO HEF calculations has recently been proposed van Hameren and Nefedov (2025). In this Letter, we will demonstrate that quite a few problems that arise when using the CF to produce MN dijet can be resolved using the HEF instead.
Cross section for the MN dijet hadroproduction within the HEF is calculated as follows:
$\sigma^{\rm(HEF)}$ $=$ $\int\frac{dx_{1}}{x_{1}}\int\limits_{{\bf q}_{T1}}\frac{1}{\pi}\,\Phi_{i}\left(x_{1},{\bf q}_{T1}^{2},\mu_{F}^{2}\right)$ (2)
$\times$ $\int\frac{dx_{2}}{x_{2}}\int\limits_{{\bf q}_{T2}}\frac{1}{\pi}\,\Phi_{j}\left(x_{2},{\bf q}_{T2}^{2},\mu_{F}^{2}\right)$
$\times$ $\hat{\sigma}_{ij}\left(x_{1,2},{\bf q}_{T1,2},\alpha_{s}(\mu_{R}^{2})\right),$
where $i,j$ denote effective gauge-invariant degrees of freedom of the EFT Lipatov (1995); Lipatov and Vyazovsky (2001)–Reggeized gluons ($R$) and quarks ($Q$). The HEF HSC $\hat{\sigma}_{ij}$ in Eq. (2) is expressed in terms of squared amplitudes with the Reggeized partons in a standard way Karpishkov et al. (2017); Nefedov and Saleev (2020). In Eq. (2), $\Phi_{i,j}$ denote UPDFs, which are responsible for the resummation of initial state radiation (ISR) effects in the HEF. We apply Kimber-Martin-Ryskin-Watt (KMRW) formula for the UPDFs Kimber et al. (2001); Watt et al. (2003) modified in Ref. Nefedov and Saleev (2020).
Introducing the variable $\bar{Y}=\alpha_{s}(\mu_{R}^{2})\,Y$, one can distinguish two regimes. When $\bar{Y}\ll 1$, the BFKL resummation of Eq. (1) is negligible and the HEF HSC can be calculated in the fixed order in $\alpha_{s}$ (“the HEF contribution”). The HEF HSC $\hat{\sigma}^{\rm(HEF)}_{ij}$ for a complete set of the LO EFT diagrams for dijet production was calculated in Ref. Nefedov et al. (2013), which is the current state of the art of the fixed-order HEF calculations.
When $\bar{Y}\sim 1$, the BFKL resummation in Eq. (1) should be taken into account (“the BFKL-improved HEF contribution”). The HSC for the MN dijet production within the BFKL formalism is factorized as follows:
$\hat{\sigma}^{\rm(BFKL)}_{ij}$ $=$ $\int\limits_{{\bf l}_{T1}}\Psi^{+}_{i}\left({\bf l}_{T1},{\bf p}_{T1},y_{1}\right)$ (3)
$\times$ $\int\limits_{{\bf l}_{T2}}\Psi^{-}_{j}\left({\bf l}_{T2},{\bf p}_{T2},y_{2}\right)\ G\left({\bf l}_{T1},{\bf l}_{T2},Y\right),$
where $\Psi^{\pm}_{i,j}$ are the IFs describing production of the jets. The IFs can be calculated in an analogous way to that explained in Refs. Ioffe et al. (2010); Kovchegov and Levin (2013), but with the Reggeized initial partons. Despite the expressions for the IFs being LO in $\alpha_{s}$, in the present Letter we go beyond the accuracy of the CF-based analysis of Refs. Colferai et al. (2010); Caporale et al. (2013) by virtue of logarithmically enhanced ISR effects taken into account by the UPDFs. One can understand this as taking a class of logarithmic corrections to the IFs of Refs. Colferai et al. (2010); Caporale et al. (2013), which are defined within the CF in these studies.
In the present study, we use the exact Green’s function $G$ of the NLO BFKL equation (1). The Green’s function with the initial condition $G\left({\bf l}_{T1},{\bf l}_{T2},Y=0\right)=\delta^{(2)}\left({\bf l}_{T1}-{\bf l}_{T2}\right)$ is constructed in Refs. Chirilli and Kovchegov (2013, 2014) as an expansion over the eigenfunctions $H_{n,\gamma}$ of the NLO kernel $K$ of Eq. (1):
$G\left({\bf l}_{T1},{\bf l}_{T2},Y\right)$ $=$ $\sum\limits_{n}\int\limits\frac{d\gamma}{2\pi i}\ \exp\left[\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi(n,\gamma)\,Y\right]$ (4)
$\times$ $H_{n,\gamma}({\bf l}_{T1},\mu_{R}^{2})\,H_{n,\gamma}^{\star}({\bf l}_{T2},\mu_{R}^{2}),$
where $\bar{\alpha}_{s}=(C_{A}/\pi)\,\alpha_{s}$. The usage of the NLO basis of the eigenfunctions, which diagonalize the kernel of Eq. (1), correctly implements the renormalization group (RG) invariance of the NLL Green’s function (4), see End Matter materials for details.
Finally, we can match the two contributions, HEF (applicable at $\bar{Y}\ll 1$) and BFKL-improved HEF (for $\bar{Y}\sim 1$), using the following matching scheme He et al. (2019):
$\hat{\sigma}_{ij}^{\rm(HEF+BFKL)}$ $=$ $\hat{\sigma}_{ij}^{\rm(HEF)}+\hat{\sigma}_{ij}^{\rm(BFKL)}-\hat{\sigma}_{ij}^{\rm(BFKL,0)},$ (5)
where $\hat{\sigma}_{ij}^{\rm(BFKL,0)}$ is the BFKL HSC in Eq. (3) calculated with the Green’s function substituted by it’s initial condition. The $\hat{\sigma}_{ij}^{\rm(BFKL,0)}$-term is the $Y\to\infty$ (Regge) asymptotics of the $\hat{\sigma}_{ij}^{\rm(HEF)}$, so it cancels the latter one for $\bar{Y}\sim 1$, and the $\hat{\sigma}_{ij}^{\rm(BFKL)}$ when $\bar{Y}\ll 1$, thus avoiding any double-counting and smoothly interpolating our predictions between the two regimes.
The renormalization and factorization scales are chosen to be equal $\mu_{R}=\mu_{F}=2^{\xi}\sqrt{|{\bf p}_{T1}|\,|{\bf p}_{T2}|}$, with $\xi=0$ for the central curves and $\xi=\pm 1$ to estimate the scale uncertainty of our predictions. The numerical integration had been performed using CUBA library Hahn (2005). We have used the modified KMRW UPDFs from Ref. Nefedov and Saleev (2020) based on parameterization of the MSTW-2008 PDFs Martin et al. (2009).
Figure 1: The $Y$ distributions of the MN jets at different energies. Several contributions are shown: HEF (solid line with yellow band), BFKL(0) (red dotted line), LO BFKL (dashed line), and NLO BFKL (dot-dashed line with blue band). The data are from CMS Collaboration Tumasyan et al. (2022).
Figure 2: The normalized $\Delta\phi$ distributions of the MN dijet at different energies and in different rapidity regions. The notations are the same as in Fig. 1. The data are from $\O$ Abachi et al. (1996) and CMS Khachatryan et al. (2016) Collaborations.
Now we are in a position to compare the results of our computations with a subset of available data on the MN dijet production at hadron colliders. Our predictions for the distributions of the rapidity difference $Y$ of the MN dijets produced in $pp$ collisions at CERN LHC at $\sqrt{S}=2.76$ TeV with $|{\bf p}_{T1,2}|>35$ GeV Chatrchyan et al. (2012) are shown in Fig. 1 (a). The HEF contribution describes the experimental data well at low $Y<2-3$ and gets subtracted by the BFKL(0) contribution starting from $Y>3-3.5$ where the transition to the BFKL regime happens. In the region of large $Y>3-3.5$, the NLO BFKL-improved HEF computation is in agreement with the data, unlike the fixed-order HEF or the LO BFKL-improved HEF, clearly showing the importance of NLL BFKL corrections. The uncertainties of the NLO BFKL-improved HEF with the Green’s function calculated by the technique of Refs. Chirilli and Kovchegov (2013, 2014) are of the same order as for the fixed-order HEF calculation, which demonstrates the reduction of scale sensitivity of the calculations in this approach. To make an unambiguous conclusion about manifestations of the BFKL effects, it is necessary to have data at higher LHC energies $\sqrt{S}=7-13$ TeV and to consider the ratio of cross sections at different energies.
Another interesting observable is the ratio $R^{\rm MN}_{\rm ex}$ of the MN dijet cross section to a certain “exclusive” dijet cross section with the veto on extra jets, designed to suppress the ISR effects Chatrchyan et al. (2012); Tumasyan et al. (2022). In the HEF approach with the KMRW UPDFs, this ISR veto can be implemented by setting the upper limit on ${\bf q}_{T1,2}$ in Eq. (2) as it was done in Ref. Nefedov et al. (2013). Our predictions at two energies, $\sqrt{S}=2.76$ and $7$ TeV, compared to the experimental data are shown in Fig. 1 (b), (c) respectively. The fixed-order HEF predictions are in satisfactory agreement with the data at low $Y<3-3.5$, while for $Y>3.5$ only the NLO BFKL-improved HEF predictions reproduce the data, whereas the LO BFKL-improved prediction gives a significantly lower value for the ratio. This again demonstrates the importance of taking into account the NLL BFKL corrections.
It had been understood for a long time in collider physics that the observable most sensitive to the QCD radiation pattern, including the BFKL effects, is the correlation of the jets in azimuthal angle. The NLO BFKL computation in the CF was found to be unstable at $\Delta\phi\to 0$ due to large $\ln\left({\bf l}_{T1,2}^{2}/\mu_{R}^{2}\right)$ in the NLO CF IFs Colferai et al. (2010). To cure this instability, the application of the Sudakov resummation through the usage of the transverse momentum dependent (TMD) factorization has been proposed Mueller et al. (2016). In our NLO BFKL-improved HEF computation we do essentially the same resummation, since the modified KMRW uPDFs are consistent with the Sudakov resummation in the TMD factorization up to NLL level as shown in Ref. Nefedov and Saleev (2020). Moreover, due to the usage of the NLO eigenfunctions of Refs. Chirilli and Kovchegov (2013, 2014) in our computation, we do not need to use any special scale-setting prescription, such as Brodsky-Lepage-Mackenzie (BLM) procedure for the renormalization scale setting Brodsky et al. (1983), which was done, e.g., in Refs. Brodsky et al. (1999); Ducloué et al. (2014); Caporale et al. (2015); Celiberto and Papa (2022).
Our predictions for the azimuthal correlation spectra at $\sqrt{S}=1.8$ and $7$ TeV and in different rapidity regions are shown in Fig. 2. The kinematic cuts of the measurement by $\O$ Abachi et al. (1996) are asymmetric, with minimum $|{\bf p}_{T}|$ of the leading(/subleading) jet being 50 GeV (/20 GeV), while the cuts for CMS data Khachatryan et al. (2016) are symmetric: $|{\bf p}_{T1,2}|>35$ GeV. At relatively small rapidities, the HEF contribution is enough to describe the data, see Fig. 2 (a), (b), (d), (e). The correct shape of the correlation spectra is evidence of the Sudakov resummation implemented in the KMRW UPDFs. At large $Y$, the HEF predicts stronger correlation between the MN jets than it is observed in the data, see Fig. 2 (c), (f). The necessary additional decorrelation comes from the radiations between the jets, which is taken into account by the BFKL resummation (1). The LO BFKL-improved HEF predicts even too strong decorrelation at both $\sqrt{S}=1.8$ TeV and $\sqrt{S}=7$ TeV, leading to an almost flat angular distribution. A good description of the data can be achieved only by using the NLL BFKL resummation, which predicts the correct shape of the MN jets correlation spectra at large rapidities (Fig. 2 (c), (f)). In such a way, the HEF with the Sudakov resummation in the UPDFs together with the BFKL resummation in the HSC describes the data on the azimuthal correlation spectra well.
To summarize, for the first time we have combined the resummation of the ISR effects provided by the HEF with the NLL BFKL resummation to describe the MN dijet hadroproduction. The HEF describes the data well at small $Y$, while at large $Y$ the NLL BFKL resummation significantly improves agreement with the data. The adopted matching scheme smoothly interpolates predictions between the two descriptions. We have used the technique of Refs. Chirilli and Kovchegov (2013, 2014) to calculate the NLL BFKL Green’s function, which reduces the renormalization scale uncertainty due to correct implementation of the RG-invariance at the NLL. The usage of the HEF with the KMRW UPDFs implementing the Sudakov resummation together with the NLO solution for the BFKL Green’s function eliminates the need for tuning of the scale choice in the MN dijet production process. Thus, evidence for the manifestations of the BFKL effects becomes more unambiguous opening prospects for their future studies at HL-LHC, CERN FCC, etc.
Acknowledgments: Authors are grateful to A. Kotikov and A. Onishchenko for valuable feedback. The work of A.C. and V.S. is supported by the BASIS Foundation for the Advancement of Theoretical Physics and Mathematics, grant #24-1-1-16-5. The work of M.N. is supported by binational Science Foundation grants #2012124 and #2021789, and by the ISF grant #910/23.
Appendix A End Matter materials.
The partonic IFs in Eq. (3) are:
$\frac{d\Psi_{R/Q}^{\pm}\left({\bf l}_{T},{\bf p}_{T},y\right)}{d^{2}{\bf p}_{T}\,dy}$ $=$ $\alpha_{s}(\mu_{R}^{2})\,\frac{C_{A/F}}{\sqrt{2}}\frac{1}{{\bf p}_{T}^{2}}$
$\times$ $\delta(1-x/x_{g/q})\ \delta^{(2)}({\bf q}_{T}\mp{\bf l}_{T}-{\bf p}_{T}).$
Due to the lack of conformal symmetry in QCD, the complete basis of orthogonal eigenfunctions $H_{n,\gamma}$ of the BFKL kernel $K$ in Eq. (1) must be constructed order-by-order in $\alpha_{s}$, ensuring the RG-invariance of the Green’s function at NLL accuracy as Chirilli and Kovchegov (2013, 2014):
$\int\limits_{{\bf q}_{T}}K\left({\bf l}_{T},{\bf q}_{T}\right)\,H_{n,\gamma}({\bf q}_{T},\mu_{R}^{2})=$ (6)
$\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi(n,\gamma)\,H_{n,\gamma}({\bf l}_{T},\mu_{R}^{2}).$
In Eq. (6), $\chi(n,\gamma)$ is the corresponding NLO BFKL characteristic function, given by:
$\chi(n,\gamma)=\chi^{(0)}(n,\gamma)+\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi^{(1)}(n,\gamma),$ (7)
where
$\chi^{(0)}(n,\gamma)$ $=$ $2\psi(1)-2\,{\rm Re}\,\psi\left(\frac{n}{2}+\gamma\right),$
$\chi^{(1)}(n,\gamma)$ $=$ $\frac{1}{2}\,\frac{\beta_{0}}{4\pi}\,\frac{\partial\chi^{(0)}(n,\gamma)}{\partial\gamma}+\frac{\delta(n,\gamma)}{4},$
the function $\delta(n,\gamma)$ is given by Eq. (27) in Ref. Kotikov and Lipatov (2000). Using the technique of Refs. Chirilli and Kovchegov (2013, 2014), the NLO eigenfunctions $H_{n,\gamma}$ (6) are constructed as:
$H_{n,\gamma}({\bf l}_{T},\mu_{R}^{2})=f^{(0)}_{n,\gamma}({\bf l}_{T})+\bar{\alpha}_{s}(\mu_{R}^{2})\,f^{(1)}_{n,\gamma}({\bf l}_{T},\mu_{R}^{2}),$ (8)
where
$f^{(0)}_{n,\gamma}({\bf l}_{T})$ $=$ $\frac{1}{\sqrt{\pi}}\,{\bf l}_{T}^{2(\gamma-1)}e^{in\phi_{l}},$
$f^{(1)}_{n,\gamma}({\bf l}_{T},\mu_{R}^{2})$ $=$ $\sum\limits_{m=0}^{\infty}\,c_{m}(n,\gamma)\ln^{m}\left(\frac{{\bf l}_{T}^{2}}{\mu_{R}^{2}}\right)f^{(0)}_{n,\gamma}({\bf l}_{T}),$
with: $c_{0}=0$, $c_{1}(n,\gamma)=\partial_{\gamma}c_{2}(n,\gamma)$, $c_{2}(n,\gamma)=\beta_{0}/\left(8\pi\,\partial_{\gamma}\ln\chi^{(0)}(n,\gamma)\right)$, and $c_{m>2}=0$ at the NLO. The integral along the contour ${\rm Re}\,\gamma=1/2$ in Eq. (4) should be understood in the sense of the principal value treating the eigenfunctions as distributions due to the pole of the $c_{2}$ at $\gamma=1/2$.
Figure 3: The (a) characteristic function (7) at $\text{Re}\,\gamma=1/2$ as function of $\text{Im}\,\gamma$ for $n=0$ and $\mu_{R}=10$ GeV and (b) intercept dependence on $\mu_{R}$. Several options are shown: LO (dashed line), NLO (dot-dashed line), and CI NLO (solid line).
The Pomeron intercept characterizing the behavior of the Green’s function (4) at $Y\to\infty$ is defined at the saddle point $\gamma$ of the appropriate characteristic function (7): $\omega=\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi(0,\gamma)$. At the LO, the latter one has only one saddle point at $\gamma=1/2$, which splits into two complex conjugate saddle points at the NLO, whose position depends on $\alpha_{s}$, see Fig. 3 (a). Our results for the characteristic function and intercept are consistent with ones presented in the recent Ref. Polizzi et al. (2025).
Figure 4: The plots of the function (10) calculated with the LL (dashed line) and CI NLL (solid line) Green’s function. In all plots except the ones showing the $\Delta\psi$-dependence, the function $g$ averaged over $\Delta\psi$ is plotted.
As it has been first observed in Ref. Salam (1998), the numerically largest NLO contribution to the characteristic function comes from (anti-)collinear poles, located at $\gamma=-n/2$ ($21+n/$ 2), which can be written as:
$\chi^{(l)}(n,\gamma)=\sum_{k=1}^{2l+1}\,\kappa_{l,k}\,d_{0}^{-k}\,\bar{\alpha}_{s}^{l}+{\cal O}(\gamma)+\{\gamma\to 1-\gamma\},$
with $\bar{\alpha}_{s}=\bar{\alpha}_{s}(\mu_{R}^{2})$, $d_{m}=d_{m}(n,\gamma)=m+\gamma+n/2$, and the coefficients $\kappa_{l,k}=\kappa_{l,k}(n)$ being defined from the pole structure of the fixed-order terms in Eq. (7). The effects of leading and sub-leading poles can be resummed Salam (1998); Sabio Vera (2005); Sabio Vera and Schwennsen (2007) to all orders in $\alpha_{s}$, the “all-poles” approximation of Refs. Sabio Vera (2005); Sabio Vera and Schwennsen (2007) for such a resummation has the form:
$\bar{\alpha}_{s}\,\Delta\chi(n,\gamma)$ $=$ $\sum_{m=0}^{\infty}\bigg{[}\kappa_{1,2}\,\bar{\alpha}_{s}-d_{m}$
$+$ $\sqrt{2\bar{\alpha}_{s}\left(\kappa_{0,1}+\kappa_{1,1}\,\bar{\alpha}_{s}\right)+\left(\kappa_{1,2}\,\bar{\alpha}_{s}-d_{m}\right)^{2}}$
-- $\sum_{l=0}^{1}\sum_{k=1}^{2l+1}\,\kappa_{l,k}\,d_{m}^{-k}\,\bar{\alpha}_{s}^{l+1}\bigg{]}+\{\gamma\to 1-\gamma\},$
where in the last line the fixed-order terms present in Eq. (7) are subtracted to avoid double-counting. Note that adding the term (A) to the characteristic function (7) is consistent with the Eq. (6) since $\Delta\chi$ is ${\cal O}(\bar{\alpha}_{s}^{3})$. In the present paper, we refer to the addition of $\Delta\chi$ to Eq. (7) as the collinear improvement (CI).
The CI NLO characteristic function has only one saddle point preserving the LO definition of the intercept at $\gamma=1/2$, see Fig. 3 (solid line). Both, NLO and CI NLO intercepts, are approximately scale independent at relatively large $\mu_{R}$ typical for the MN dijet production and are much smaller than the LO intercept.
It is convenient, in particular for numerical interpolation purposes, to work with the function
$g\left({\bf l}_{T1},{\bf l}_{T2},Y\right)$ $=$ $\frac{|{\bf l}_{T1}||{\bf l}_{T2}|\,G\left({\bf l}_{T1},{\bf l}_{T2},Y\right)}{G_{\rm D}\left(|{\bf l}_{T1}|/|{\bf l}_{T2}|,Y\right)},$ (10)
which is the exact Green’s function (4) divided by the diffusion approximation valid at $\bar{Y}\sim 1$, in LL approximation:
$G_{\rm D}\left(|{\bf l}_{T1}|/|{\bf l}_{T2}|,Y\right)$ $=$ $\frac{1}{2\pi^{2}}\sqrt{\frac{\pi}{D\,\bar{Y}}}$
$\times$ $\exp\left[\omega_{0}\,\bar{Y}-\frac{1}{D\,\bar{Y}}\,\ln^{2}\left(\frac{|{\bf l}_{T1}|}{|{\bf l}_{T2}|}\right)\right],$
here $\omega_{0}=4\ln(2)$ and $D=14\zeta(3)$. The dependence of the function (10) on the momenta ${\bf l}_{T1,2}$ can be conveniently parameterized with the help of variables: $L=\ln\left(|{\bf l}_{T1}|/|{\bf l}_{T2}|\right)$, $\Delta L=\ln\left(|\Delta{\bf l}_{T}|/\mu_{R}\right)$, where $\Delta{\bf l}_{T}={\bf l}_{T1}-{\bf l}_{T2}$, and $\Delta\psi$ being the azimuthal angle between ${\bf l}_{T1,2}$.
Several plots of the function (10) are shown in Fig. 4. At relatively small $\bar{Y}\lesssim 0.5$, the behavior of the function $g$ w.r.t. $L$ and $\Delta\psi$ approximates the initial condition of the Green’s function (4): $g\sim\delta(L)\,\delta(\Delta\psi)$. With rapidity increasing up to $\bar{Y}\sim 1.5$, the peak spreads and the LL $g$ function approaches unity due to the validity of the diffusion approximation in this region. The CI NLL Green’s function does not show any unphysical oscillations or negative values, as evident from Fig. 4. We also point out that the effect of CI on the Green’s function is numerically small, consistently with observations of Refs. Colferai et al. (2010); Ducloue et al. (2013); Ducloué et al. (2014) and the NLO approximation (7) is actually sufficient for the description of the data. The dependence on $\Delta L$ of the NLL $g$ is rather small, so we do not vary it in Fig. 4 setting $\Delta L=0$, i.e. $\mu_{R}=|\Delta{\bf l}_{T}|$. For $\bar{Y}\sim 1$, the NLL $g$-function is almost flat, but a few times lower than the LL one due to the smaller intercept value (Fig. 3 (b)). Thus, the full NLL Green’s function ($G$) grows slower with increasing of $\bar{Y}$ than the LL one.
In this way, the implementation of the NLO basis of the eigenfunctions together with the CI characteristic function provides the RG-invariant NLL Green’s function without any pathologies.